1. If you draw a random sample of size 80 from an infinitely large population with a variance of 6, what is the standard error of the mean? 2. Now suppose the population from question 1 is of size 150, what is the standard error of the mean?


1
Expert's answer
2021-07-08T15:19:48-0400

1.


"\frac{\sigma}{\sqrt{n}}=\frac{6}{8.9}=0.67"

2.


"\frac{6}{12.2}=0.5"